OpenStudy (anonymous):

Solve the system by the method of your choice x=2y-9 -3x+6y=27

OpenStudy (ajprincess):

substitute the first equation in the 2nd. then u will get -3(2y-9)+6y=27 -6y+27+6y=27 since y cancel out you will not have any solutions.